# TACO: Enhancing Multimodal In-context Learning via Task Mapping-Guided Sequence Configuration

### v1

@@ Abstract
Multimodal in-context learning (ICL) has emerged as a key mechanism for harnessing the capabilities of large vision–language models (LVLMs). However, its effectiveness remains highly sensitive to the quality of input in-context sequences, particularly for tasks involving complex reasoning or open-ended generation. A major limitation is our limited understanding of how LVLMs actually exploit these sequences during inference. To bridge this gap, we systematically interpret multimodal ICL through the lens of task mapping, which reveals how local and global relationships within and among demonstrations guide model reasoning. Building on this insight, we present TACO, a lightweight transformer-based model equipped with task-aware attention that dynamically configures in-context sequences. By injecting task-mapping signals into the autoregressive decoding process, TACO creates a bidirectional synergy between sequence construction and task reasoning. Experiments on five LVLMs and nine datasets demonstrate that TACO consistently surpasses baselines across diverse ICL tasks. These results position task mapping as a valuable perspective for interpreting and improving multimodal ICL.
@@ 1 Introduction
In-context learning (ICL) is a paradigm in which models make predictions by conditioning on a sequence of input–output demonstrations, without updating their parameters. This approach enables models to rapidly adapt to new tasks using only a few examples provided at inference time. Initially, ICL gained significant traction in the domain of large language models (LLMs), where it has demonstrated impressive performance across a wide range of tasks Olsson et al. (2022); Garg et al. (2023). More recently, the concept has been extended to the multimodal setting, giving rise to multimodal ICL, where large vision–language models (LVLMs) learn from interleaved image–text sequences and support complex multi-image reasoning. This capability has become a cornerstone of modern LVLMs Alayrac et al. (2022); Chen et al. (2025); Bai et al. (2025), enabling more flexible and generalizable multimodal understanding.
Despite significant progress in multimodal ICL, configuring effective input sequences remains an open challenge. A standard ICL sequence consists of an instruction, a set of in-context demonstrations (ICDs), and a query sample (see Figure 1). Prior studies have shown that even small changes to these sequences can substantially alter LVLM predictions Schwettmann et al. (2023); Zhou et al. (2024); Li et al. (2025). These findings highlight the need for robust configuration strategies. However, as multimodal ICL involves diverse cross-modal interactions, our mechanistic understanding is still limited. As a result, existing methods depend on hand-crafted metrics to assess each ICD’s contribution to LVLM reasoning Iter et al. (2023); Fan et al. (2024). Rather than relying on specific metrics, we propose a more effective model-centric alternative. To this end, we address two core research questions:
@@ 2 Task Mapping in Multimodal ICL
2.1 Identifying Task Mapping in VL ICL
Notations. In this work, we mainly focus on ICL for image-to-text tasks, where ICL sequences are organized in an interleaved image-text format. Toward a unified template for various tasks, we reformat ICDs as triplets $(I,Q,R)$, where $I$ is an image, $Q$ is a task-specific text query and $R$ is the ground-truth response. The query sample is denoted as $(\hat{I},\hat{Q})$. Formally, ICL can be represented as:
$\hat{R}\leftarrow\mathcal{M}(S^{n})=\mathcal{M}(Inst;\underbrace{(I_{1},Q_{1},R_{1}),...,(I_{n},Q_{n},R_{n})}_{n\times ICDs};(\hat{I},\hat{Q})),$
where $\mathcal{M}$ is a pretrained LVLM, $S^{n}$ is an ICL sequence consists of an instruction $Inst$, $n$-shot ICDs and a query sample, as illustrated in Figure 1.
Definition. We define task mapping as a model-learnable inferential process that transforms input modalities into their outputs within the LVLM’s latent space, capturing both local and global relationships in ICL. Consequently, each ICD $(I_{i},Q_{i},R_{i})$ possesses a local task mapping:
@@ 2.1 Identifying Task Mapping in VL ICL
specifying how its image and query jointly map to the response. Then LVLM’s generation on the target query sample $(\hat{I},\hat{Q})$ can be viewed as a global task mapping:
@@
Visualization. We employ a specific-mapping task, HatefulMemes Kiela et al. (2020), to reveal task mapping. Here, each local mapping $f_{i}$ is defined by a binary classification: given a meme image with its caption, determine whether it contains harmful content and output "yes" or "no." (Figure 1(a)) We use the validation split as our query set and sample $n$ ICDs from the training split using Random Sampling (RS) with a normal distribution to configure the ICL sequences. To highlight $f_{i}$, we create two setups:
@@ 2.2 Task Mapping is Key to Multimodal ICL
Results. Figure 2 shows the LVLMs’ ICL performance on the same sequences under different settings. The findings are as follows:
Figure 2: Results on Hatefulmemes under various settings. "+" denotes combined settings.
Better capturing task mapping consistently improves performance. As shown in Figure 2(a), across all shot counts, EM > Standard > HM in a clear descending order. This aligns with our observations from in-context lens. In Figure 2(b), removing instructions, which serve as higher-level guidance enabling LVLMs to more deeply capture and utilize $f_{i}$, generally lowers performance. Yet “EM w/o $Inst$” still surpasses Standard.
@@ 2.3 ICL Needs Cohesive Global Mapping
Setups. Three sequence configuration methods are evaluated: Random Sampling (RS), similarity-based retrieval, and an idealized Oracle. In similarity-based retrieval, ICDs are selected by CLIP cosine similarity, using either image only alignment (I2I) or joint image query alignment (IQ2IQ). The Oracle method greedily chooses each ICD to maximize the log likelihood of generating the ground-truth response while accounting for the cumulative influence of prior demonstrations (computational details in Appendix C.3).
@@ 3 The Proposed Method: TACO
Input Embedding. Let $x_{i}$ denote $i$-th ICD token $(I_{i},Q_{i},R_{i})$ and $\hat{x}$ denote the query sample $(\hat{I},\hat{Q})$. In each input sequence, $\hat{x}$ is placed ahead of all $x_{i}$. To align with the autoregressive generation process, we use two special tokens, $[BOS]$ and $[EOS]$, to mark the beginning and end of the input sequence during training. These tokens are added to TACO’s vocabulary. We also introduce a $[TASK]$ token into the vocabulary and concatenate it with $\hat{x}$ in the input sequence. It acts as a semantic anchor for task mapping. Therefore, for a given $S^{N}$, we reconstruct it as a token sequence $([BOS],[TASK]+\hat{x},x_{1},...,x_{N},[EOS]\})$. To filter and balance multimodal features for better mapping construction, we employ a binary fusion module to generate the embedding $e_{i}$ for $x_{i}$:
@@
where $\mathrm{MLP}^{(l)}$: $\mathbb{R}^{2d}\rightarrow\mathbb{R}^{d}$ is a layer-specific network producing a scalar weight $t_{i}^{l}\in[0,1]$ and $\sigma$ is the sigmoid function. This weight reflects the degree to which each token contributes to the cohesive task mapping, dynamically adapting TACO’s attention to emphasize semantically salient features. It modulates attention logits through a task-aware mask $M^{(l)}$. For intra-ICD tokens, the mask scales pairwise cosine similarities by $log(t_{i}^{(l)})$. For query-ICD tokens, a learnable coefficient $\alpha$ allows $\hat{e}$ to guide attention throughout the sequence. The mask is computed as follows for position $(i,j)$:
$M_{ij}^{(l)}=\begin{cases}\frac{\mathrm{sim}(e_{i},\,e_{j})}{\sqrt{d}}\;\cdot\;\log\bigl{(}t_{i}^{(l)}\bigr{)},&j\leq i\text{ and }i,j\in I_{idx},\\[4.0pt] \frac{\alpha\mathrm{sim}(\hat{e},\,e_{j})}{\sqrt{d}}\;\cdot\;\log\bigl{(}t_{1}^{(l)}\bigr{)},&i=1\text{ and }j\in I_{idx},\\[4.0pt] -\infty,&\text{otherwise}.\end{cases}$
Here, the first case emphasizes interactions between local task mappings and the second case enable deep task mapping cohesion. The last case preserves the autoregressive nature. The mask $M^{(l)}$ is integrated into standard attention, forming task-aware attention (TaAttn), as follows:
@@
where $\mathcal{U}$ is a uniform distribution. Minimizing this KL divergence prompt a sharper representation of task-mapping. The total training objective combines the standard cross-entropy loss for sequence generation, sparsity regularization, and L2-norm constraint on $TG$ to prevent overfitting:
@@ 4.1 Training Data Construction and Models
Following standard multimodal ICL evaluation practices Awadalla et al. (2023), we select six high-quality datasets across three core VL tasks: VQAv2, VizWiz Gurari et al. (2018), and OK-VQA Marino et al. (2019) for open-ended VQA; Flickr30K Young et al. (2014) and MSCOCO Lin et al. (2014) for captioning; and HatefulMemes for classification. To further assess TACO’s abilities in generalized-mapping tasks, we create a mixed-task dataset Hybrid, by sampling 5,000 instances from each above dataset’s training set, with validation samples drawn proportionally from their validation sets. We also adopt two challenging image-to-text tasks from the latest multimodal ICL benchmark, VL-ICL Zong et al. (2024): Fast Open-Ended MiniImageNet [^1] We apply fine-grained adjustments to Fast to increase its difficulty; see Appendix E.1 for details. (Fast) and CLEVR.
@@
Our experiments evaluate four advanced open-source LVLMs: OpenFlamingov2-9B, Idefics2-8B, InternVL2.5-8B Chen et al. (2025), and Qwen2.5VL-7B Bai et al. (2025), as well as a closed-source model, GPT-4V OpenAI et al. (2024). Detailed descriptions of the datasets and LVLMs are provided in Appendix E.1.
@@ 4.2 Baselines and Implementation Details
We adopt RS and two similarity-based retrieval methods introduced in §2.3 as baselines, along with three previous SOTA configuration methods: IQPR Li et al. (2024), DEmO Guo et al. (2024), and Lever-LM Yang et al. (2024). Lever-LM is a tiny language model with several standard decoder blocks that performs automatic $S^{n}$ configuration. As it also requires model training, we treat it as the primary baseline. For a fair comparison, we set Lever-LM’s depth to four layers. Details of the baselines are provided in Appendix E.3.
@@
Table 2: Results of the ablation study on task mapping augmentation of TACO. Specifically, (a)-(d) correspond to diverse task-aware attention construction, (e)-(h) to diverse $TG$ initialization.
Figure 5: Results of TACO with and without task-aware attention under different $N$-$n$ settings across three datasets, where $N$ is the training sequence shot and $n$ is the generation sequence shot.
@@ 4.3 Main Results
Table 1 summarizes the average ICL performance across five LVLMs under different ICL sequence configuration methods. TACO consistently outperforms all baselines across all nine datasets, highlighting its robustness and effectiveness in fully leveraging the potential of LVLMs for diverse multimodal ICL scenarios. Notably, TACO delivers particularly strong results in generalized-mapping tasks, achieving an average improvement of 3.26% in VQA tasks, with the second highest gain of 3.61% observed on Hybrid. These results demonstrate that strengthening task mapping enhances the autoregressive generation process of language models, equipping them with a broader understanding and enabling the construction of more precise cohesive task mappings. In Appendix E.4, we present more evaluations of how ICL sequence configuration affects LVLM using per-model data and include efficiency analyses of TACO to show its low computational overhead.
@@ 5 Ablation Study and Discussions
Task mapping consistently enhances multimodal ICL. Figure 5 shows that across all $N$-$n$ combinations, task-aware attention always improves performance, highlighting the value of focusing ICL sequences on task mapping.
@@ 6 Related Works
Interpreting ICL. The mechanisms of ICL are crucial to better employing it Gao et al. (2021); Dong et al. (2024); Li (2025). Min et al. (2022) attribute ICL’s success to explicit information in ICDs like label space and input distribution, while Zhou et al. (2023) emphasize the importance of input-output mappings. To find a unified solution, Wei et al. (2023) and Pan et al. (2023) disentangle ICL into Task Recognition and Task Learning. Zhao et al. (2024) further propose a two-dimensional coordinate system to explain ICL behavior via two orthogonal variables: similarity in ICDs and LLMs’ ability to recognize tasks. However, these studies are often confined to specific-mapping tasks with small label spaces and struggle to address complex multimodal scenarios.

### v2

@@ Abstract
Multimodal in-context learning (ICL) has emerged as a key mechanism for harnessing the capabilities of large vision–language models (LVLMs). However, its effectiveness remains highly sensitive to the quality of input ICL sequences, particularly for tasks involving complex reasoning or open-ended generation. A major limitation is our limited understanding of how LVLMs actually exploit these sequences during inference. To bridge this gap, we systematically interpret multimodal ICL through the lens of task mapping, which reveals how local and global relationships within and among demonstrations guide model reasoning. Building on this insight, we present TACO, a lightweight transformer-based model equipped with task-aware attention that dynamically configures ICL sequences. By injecting task-mapping signals into the autoregressive decoding process, TACO creates a bidirectional synergy between sequence construction and task reasoning. Experiments on five LVLMs and nine datasets demonstrate that TACO consistently surpasses baselines across diverse ICL tasks. These results position task mapping as a novel and valuable perspective for interpreting and improving multimodal ICL.
@@ 1 Introduction
In-context learning (ICL) is a paradigm in which models make predictions by conditioning on a sequence of input–output demonstrations, without updating their parameters. This approach enables models to rapidly adapt to new tasks using only a few examples provided at inference time. Initially, ICL gained significant traction in the domain of large language models (LLMs), where it has demonstrated impressive performance across a wide range of tasks Olsson et al. (2022); Garg et al. (2023). More recently, the concept has been extended to the multimodal setting, giving rise to multimodal ICL, where large vision–language models (LVLMs) learn from interleaved image–text sequences and support complex multi-image reasoning. This capability has become a cornerstone of modern LVLMs Alayrac et al. (2022); Chen et al. (2024c); Bai et al. (2025), enabling more flexible and generalizable multimodal understanding.
Despite significant progress in multimodal ICL, configuring effective input sequences remains an open challenge. A standard ICL sequence consists of an instruction, a set of in-context demonstrations (ICDs), and a query sample (see Figure 1). Prior studies have shown that even small changes to these sequences can substantially alter LVLM predictions Schwettmann et al. (2023); Zhou et al. (2024); Li et al. (2025c). These findings highlight the need for robust configuration strategies. However, as multimodal ICL involves diverse cross-modal interactions, our mechanistic understanding is still limited. As a result, existing methods depend on hand-crafted metrics to assess each ICD’s contribution to LVLM reasoning Iter et al. (2023); Fan et al. (2024). Rather than relying on specific metrics, we propose a more effective model-centric alternative. Specifically, we address two research questions:
@@
Our main contributions can be summarized in three-fold:
  • To fill the gap in research on multimodal ICL mechanisms in LVLMs, we propose a task mapping framework that systematically captures task distributions across ICDs within an ICL sequence. This framework not only offers a unified view of multimodal ICL under diverse scenarios but also sheds light on the internal behavior of LVLMs during ICL.
  • Within this task-mapping framework, we propose TACO, a lightweight transformer-based model designed to adaptively retrieve and arrange ICDs from a dataset, yielding optimal ICL sequences for a target LVLM. Evaluations on five LVLMs and nine benchmarks show that TACO achieves superior performance over prior configuration strategies.
  • We carry out an extensive analysis and ablation study of TACO, isolating the role of each component and design decision. The results provide insights into TACO’s underlying mechanisms and further demonstrate the effectiveness of our task mapping framework for ICL enhancements and applications.
@@ 2 Task Mapping in Multimodal ICL
2.1 Identifying Task Mapping in multimodal ICL
[…]
$),\hat{R}\leftarrow\mathcal{M}(S^{n})=\mathcal{M}(Inst;\underbrace{(I_{1},Q_{1},R_{1}),...,(I_{n},Q_{n},R_{n})}_{n\times ICDs};(\hat{I},\hat{Q}$)),
where $\mathcal{M}$ is a pretrained LVLM, $S^{n}$ is an ICL sequence consisting of an instruction $Inst$, $n$-shot ICDs and a query sample, as shown in Figure 1.
Task Mapping Definition. We define task mapping as a model-learnable inferential process that transforms input modalities into their outputs within the LVLM’s latent space, capturing both local and global relationships in ICL. Each ICD $(I_{i},Q_{i},R_{i})$ possesses a local task mappingff:
@@ 2.1 Identifying Task Mapping in multimodal ICL
[…]
@@
Visualization. We employ a specific‐mapping task, HatefulMemes Kiela et al. (2020), to reveal task mapping. Here, each local mapping $f_{i}$ is defined by a binary classification: given a meme image with its caption, determine whether it contains harmful content and output "yes" or "no." (Figure 1(a)) We use the validation split as our query set and sample $n$ ICDs from the training split using Random Sampling (RS) with a normal distribution to configure the ICL sequences. To highlight $f_{i}$, we create two setups:
@@ 2.2 Task Mapping is Key to Multimodal ICL
Results. Figure 2 shows the LVLM’s ICL performance on the same sequences under different settings. The findings are as follows:
Figure 2: Results on HatefulMemes under various settings. "+" denotes combining two settings.
Better capturing task mapping consistently improves performance. As shown in Figure 2(a), across all shot counts, EM > Standard > HM in a clear descending order. This aligns with our observations from in-context lens. In Figure 2(b), removing instructions, which serve as higher-level guidance enabling LVLMs to more deeply capture and utilize $f_{i}$, generally lowers performance. Yet “EM w/o InstInst” still surpasses Standard.
@@ 2.3 ICL Needs Cohesive Global Mapping
Setups. Three sequence configuration methods are evaluated: Random Sampling (RS), similarity-based retrieval, and an idealized Oracle. In similarity-based retrieval, ICDs are selected by CLIP cosine similarity, using either image-only alignment (I2I) or joint image query alignment (IQ2IQ). The Oracle method greedily chooses each ICD to maximize the log likelihood of generating the ground-truth response while accounting for the cumulative influence of prior demonstrations (computational details in Appendix C.3).
@@ 3 The Proposed Method: TACO
[…]
@@
where $\mathrm{MLP}^{(l)}$: $\mathbb{R}^{2d}\rightarrow\mathbb{R}^{d}$ is a layer-specific network producing a scalar weight $t_{i}^{l}\in[0,1]$ and $\sigma$ is the sigmoid function. This weight reflects the degree to which each token contributes to the cohesive task mapping, dynamically adapting TACO’s attention to emphasize semantically salient features. It modulates attention logits through a task-aware mask $M^{(l)}$. For intra-ICD tokens, the mask scales pairwise cosine similarities by $\log(t_{i}^{(l)})$. For query-ICD tokens, a learnable coefficient $\alpha$ allows $\hat{e}$ to guide attention throughout the sequence. The mask is computed as follows for position $(i,j)$:
$M_{ij}^{(l)}=\begin{cases}\frac{\mathrm{sim}(e_{i},\,e_{j})}{\sqrt{d}}\;\cdot\;\bigl(-\log(t_{i}^{(l)})\bigr),&j\leq i\text{ and }i,j\in I_{idx},\\[4.0pt] \frac{\alpha\mathrm{sim}(\hat{e},\,e_{j})}{\sqrt{d}}\;\cdot\;\bigl(-\log(t_{1}^{(l)})\bigr),&i=1\text{ and }j\in I_{idx},\\[4.0pt] -\infty,&\text{otherwise}.\end{cases}$
Here, the first case emphasizes interactions between local task mappings, and the second case enables deep task mapping cohesion. The last case preserves the autoregressive nature. The mask $M^{(l)}$ is integrated into standard attention, forming task-aware attention (TaAttn), as follows:
@@
where $\mathcal{U}$ is a uniform distribution. Minimizing this KL divergence prompts a sharper representation of task-mapping. The total training objective combines the standard cross-entropy loss for sequence generation, sparsity regularization, and L2-norm constraint on $TG$ to prevent overfitting:
@@ 4.1 Training Data Construction and Models
Following standard multimodal ICL evaluation practices Awadalla et al. (2023), we select six high-quality datasets across three core VL tasks: VQAv2, VizWiz Gurari et al. (2018), and OK-VQA Marino et al. (2019) for open-ended VQA; Flickr30K Young et al. (2014) and MSCOCO Lin et al. (2014) for captioning; and HatefulMemes for classification. To further assess TACO’s abilities in generalized-mapping tasks, we create a mixed-task dataset Hybrid, by sampling 5,000 instances from the training set from each above dataset, with validation samples drawn proportionally from their validation sets. We also adopt two challenging image-to-text tasks from the latest multimodal ICL benchmark, VL-ICL Zong et al. (2024): Fast Open-Ended MiniImageNet [^1] We apply fine-grained adjustments to Fast to increase its difficulty; see Appendix E.1 for details. (Fast) and CLEVR.
@@
Our experiments evaluate four advanced open-source LVLMs: OpenFlamingov2-9B, Idefics2-8B, InternVL2.5-8B Chen et al. (2024c), and Qwen2.5VL-7B Bai et al. (2025), as well as a closed-source model, GPT-4V OpenAI (2023). Detailed descriptions of the datasets and LVLMs are provided in Appendix E.1.
@@ 4.2 Baselines and Implementation Details
We adopt RS and two similarity-based retrieval methods introduced in §2.3 as baselines, along with three previous SOTA configuration methods: IQPR Li et al. (2024b), DEmO Guo et al. (2024), and Lever-LM Yang et al. (2024). Lever-LM is a tiny language model with several standard decoder blocks that performs automatic $S^{n}$ configuration. As it also requires model training, we treat it as the primary baseline. For a fair comparison, we set Lever-LM’s depth to four layers. Details of the baselines are provided in Appendix E.3.
@@
Table 2: Results of the ablation study on task mapping augmentation of TACO. Specifically, (a)-(d) correspond to diverse task-aware attention construction, (e)-(h) to diverse TGTG initialization.
Figure 5: Results of TACO with and without task-aware attention under different NN-$n$ settings across three datasets, where $N$ is the training sequence shot and $n$ is the generation sequence shot.
@@ 4.3 Main Results
Table 1 summarizes the average performance of ICL in five LVLMs using different methods of configuring the ICL sequence. TACO consistently outperforms all baselines across all nine datasets, highlighting its robustness and effectiveness in fully leveraging the potential of LVLMs for diverse multimodal ICL scenarios. Notably, TACO delivers particularly strong results in generalized-mapping tasks, achieving an average improvement of 3.26% in VQA tasks, with the second highest gain of 3.61% observed on Hybrid. These results demonstrate that strengthening task mapping enhances the autoregressive generation process of language models, equipping them with a broader understanding and enabling the construction of more precise cohesive task mappings. In Appendix E.4, we present more evaluations of how ICL sequence configuration affects LVLM using per-model data and include efficiency analyses of TACO to show its low computational overhead.
@@ 5 Ablation Study and Discussions
Method VQAv2 VizWiz OK-VQA Flickr30K MSCOCO HatefulMemes Hybrid
TACO (RS) 62.38 47.69 54.47 98.31 115.83 76.49 38.62
TACO (I2I) 61.95 47.28 53.86 98.74 117.25 77.46 36.90
TACO (IQ2IQ) 64.37 50.18 59.23 99.17 118.68 78.93 41.05
TACO (Oracle) 66.75 52.07 61.54 99.62 119.47 80.59 45.22
@@
Task mapping consistently enhances multimodal ICL. Figure 5 shows that across all NN-$n$ combinations, task-aware attention always improves performance, highlighting the value of focusing ICL sequences on task mapping.
@@
Next, we investigate the construction of training sequences for TACO. In the main experiments, we use Oracle that approximates an LVLM’s optimal multimodal ICL. Training TACO on sequences generated by Oracle enables it to learn how to configure ICL inputs with correct task mappings, especially at the global level across the demonstration set. To test the robustness of this learning process, we replace Oracle with similarity-based methods and examine how this change affects TACO’s accuracy in inferring task mappings.
As TACO is training-based, one of the most important aspects is high-quality data. Table 3 demonstrates that using Oracle as the method to construct the training data is optimal. Since our approach leverages TACO to capture how LVLMs understand and utilize task mapping, the training data that best reflects the internal mechanisms of the LVLM is most effective. Moreover, it can be observed that TACO, when combined with RS, I2I, and IQ2IQ, brings significant performance improvements over using RS, I2I, and IQ2IQ in isolation. This indicates that our method can mitigate the inherent limitations of retrieval strategies through training, further enhancing the practical value of TACO.
@@ 6 Related Works
Interpreting ICL. The mechanisms of ICL are crucial to better employing it Gao et al. (2021); Dong et al. (2024); Li (2025). Min et al. (2022) attribute ICL’s success to explicit information in ICDs like label space and input distribution, while Zhou et al. (2023) emphasize the importance of input-output mappings. To find a unified solution, Wei et al. (2023) and Pan et al. (2023) disentangle ICL into Task Recognition and Task Learning. Zhao et al. (2024) further propose a two-dimensional coordinate system to explain ICL behavior via two orthogonal variables: similarity in ICDs and LLMs’ ability to recognize tasks. However, these studies are often confined to specific-mapping tasks with small label spaces and struggle to address complex multimodal scenarios. We present related work on configuring ICL sequences in Appendix A.